### v1

@@ Abstract
We introduce HeadZoom, a hands-free interaction technique for navigating two-dimensional visual content using head movements. The system enables fluid zooming and panning by only using real-time head tracking. It supports natural control in applications such as map exploration, radiograph inspection, and image browsing, particularly where physical interaction is limited. We evaluated HeadZoom in a within-subjects user study comparing three interaction techniques—Static, Tilt Zoom, and Parallel Zoom—across spatial, error, and subjective metrics. Results show that Parallel Zoom significantly reduced total head movement compared to Static and Tilt modes. Users reported significantly lower perceived exertion for Parallel Zoom, confirming its suitability for prolonged or precision-based tasks. By minimizing movement demands while maintaining task effectiveness, HeadZoom advances the design of head-based 2D interaction in VR and creates new opportunities for immersive, accessible, and hands-free systems for image exploration.
@@
Mixed Reality (MR) is moving beyond entertainment and into real-world domains, such as surgery, fieldwork, manufacturing, and education, where users’ hands are often occupied, gloved, sterilized, or constrained. In these contexts, reaching for a controller or touching a surface is inconvenient, impractical, unsafe, or impossible.
As MR technology advances, the demand for innovative hands-free interaction methods continues to grow, particularly in applications involving exploring and manipulating two-dimensional (2D) visual content such as maps, radiographs, and photographs. Traditional input devices, such as hand-held controllers or touchscreens, can be cumbersome or impractical in scenarios requiring hands-free interaction. To address these limitations, head-based interaction offers a promising alternative, enabling users to zoom in and pan across images using natural head movements.
HeadZoom introduces a novel framework for head-controlled zooming and panning of 2D content. The system uses precise head motion tracking to provide smooth and responsive control. This makes it particularly suitable for applications where precision and efficiency are critical, such as navigating geographical maps, examining medical radiographs, or exploring personal or professional photographs. This paper presents the development and evaluation of HeadZoom, with the following key contributions:
@@
User-Centered Design: An interface that reduces cognitive and physical effort, with efficient interaction in extended use.
Extensive Evaluation: Comparing HeadZoom with existing interaction paradigms, demonstrating its effectiveness in exploring maps, analyzing radiographs, and browsing photographs.
By redefining how head motion is used for interaction, HeadZoom provides a powerful alternative for interacting with two-dimensional visual content, supporting advances in areas such as geographic information systems, medical imaging, and digital photography.
In the remainder of this paper, we first review related work and clarify the research gap that HeadZoom addresses and the novel contribution of the work. Next, we describe the implementation of HeadZoom, before presenting a user study to evaluate the technique. Section 7 contains the user study results discussed in Section 8. Finally, we close with conclusions and future research directions.
@@ 2 Related Work
Navigating visual content often requires panning (moving the viewport) and zooming (adjusting scale). Zooming and positioning techniques are critical to improving navigation efficiency in different applications. Many modern software tools, such as image editors or maps on the Web, provide interfaces for zooming- and panning-based navigation of visual content. This section reviews existing research in these areas and highlights the role of head-based interaction in addressing current challenges.
@@ 2.1 Zooming and Panning Interfaces
Zooming and panning interfaces have been proposed for the 2D desktop GUI. You et al. [57] conducted a usability evaluation of the zoom and pan functions in web-based map interfaces, comparing different layouts of the zoom and pan buttons. Gutwin and Fedak [17] variants of panning and zooming interfaces to fit large 2D GUIs to smaller screens. Igarashi and Hinckley [23] proposed automatically controlling the zoom level proportional to the scrolling speed to help navigate large documents in a 2D GUI. A usability study showed that users can navigate through a large document efficiently and smoothly without getting disoriented, compared to manual control, which suffers from fast visual flow when scrolling too quickly while zoomed in. Cockburn et al. [13] provide a comprehensive review of early research on interfaces for navigating large documents in 2D desktop GUI, categorizing them into Overview+Detail, Zooming, Focus+Context, and Cue-based techniques. Standard input methods for zooming and panning in 2D GUI include mouse, touch gestures, and shortcut keys [57]. In addition to these popular techniques, researchers also investigated using other input methods, such as 6DOF controllers [45] or gaze input [51, 19].
The popularization of mobile devices with small screens further emphasized the need to explore zooming and panning techniques [17]. Although the pinch-to-zoom method became widely adopted with the arrival of modern touch screen-based mobile devices [1], its clear limitation of occupying both hands motivated many researchers to focus on providing a one-handed solution to control zooming and panning on mobile devices [33]. To overcome such limitations, researchers have developed various touch gestures [14, 27] to replace or complement the standard pinch-to-zoom gesture. At the same time, others also explored using additional detectors, such as pressure sensors combined with a touch screen to enable natural zooming and scrolling [35, 53], or using a gyro sensor or buttons built into the device [16].
Zooming and panning interfaces have also been explored in AR and VR applications. Typically, handheld controllers [11] or hand gestures and gaze input [47, 42, 34] are used to control zoom and pan in VR / AR applications, while some adopt touch-screen-based interaction [37, 10]. Lee et al. [28] proposed an automatic control of the zoom level based on the distance between the user’s view and the target object, increasing the zoom level as the AR camera approaches the tracking target, enabling users to have a closer look of the content with less tracking failure. While they linked the zoom control to the movement of the user’s view, the users held the AR camera in their hands. In comparison, our work investigates how head movement, as one of the hands-free input methods, could be useful for controlling zooming and panning operations in exploring large high-resolution images in VR.
@@ 2.2 Hands-Free Input Methods
Eye Tracking. Eye tracking technology has been actively researched in the context of human-computer interaction [32, 49]. Gaze direction is a natural way to control an interface, as the eyes instinctively focus on areas of interest, and has also been actively investigated in VR [44, 42] and AR [26, 39] interfaces. Modern eye trackers using infrared corneal reflections achieve high accuracy [32, 49]. A key challenge, however, is the Midas touch problem, which makes it difficult to distinguish intentional interactions from normal eye movements [25, 32]. Although some modern VR/AR HMDs integrate eye tracking technology, the technology is mostly supported on high-end devices. It has not reached the status of being widely adopted as a standard feature.
Head-Based Interaction. Head tracking has been studied to control computers using natural head movements [22, 52]. Early work in human-computer interaction identified head movement as a simple and effective pointing method, often more accurate than voice or brain-computer interfaces [49]. It is also considered to be more affordable compared to eye tracking technology.
Head-based input is especially useful in assistive technology, allowing users with limited limb mobility to move a cursor or make selections by moving their head, using a webcam or infrared sensor. Systems such as Camera Mouse [5, 38] and Enable Viacam (eViacam)111https://eviacam.crea-si.com track the face of a user through a webcam, translating head movements into pointer control [8]. These systems have been used successfully by individuals with high spinal cord injuries, as head and eye movements remain intact [5, 49].
Head tracking is also a key enabler technology for inducing a sense of presence in VR [48, 55]. Modern VR/AR HMDs typically integrate head tracking technology to use head orientation and position to control the user’s viewpoint. Harrison and Dey’s Lean and Zoom system [20] detected users leaning toward or away from the screen, automatically magnifying the content. Users reported that this head-driven zooming felt natural and fluid, making the text easier to read and reducing discomfort.
@@ 3 The HeadZoom Prototype
HeadZoom is a hands-free interaction framework that provides precise, controller-free zooming and panning in VR. As shown in Figure 2, the system continuously captures the HMD position and orientation, smoothing through Kalman filters to suppress jitter while preserving responsiveness. A virtual image plane is placed in front of the user, and raycasting from the filtered head pose determines the target location for interaction. This plane-based mapping enables natural, spatially consistent simultaneous panning and zooming using only head motion. The HMD’s position vector provides real-time 3D coordinates of the user’s head in space. This vector is used to compute the user’s distance from the image plane, which modulates the zoom level—closer positions result in a zoom-in effect, while retreating causes zoom-out. The forward vector represents the HMD directional orientation, specifically the normalized vector pointing where the user is facing. This vector is crucial for raycasting: It projects a ray from the user’s current position toward the image plane. The intersection of this ray with the plane determines the panning target, allowing the system to infer where in the image the person is attending to. Both vectors are filtered to reduce noise before being passed to the technique controller.
As illustrated in Figure 1, HeadZoom supports three head-driven interaction techniques that vary in how motion is mapped to image behavior. In the Static condition, zooming is achieved by moving closer to or farther from a fixed image in space, similar to real-world inspection. Panning is not actively supported; the image remains stationary regardless of head rotation.
In the Parallel condition, head translations toward the image plane modulate zoom level. At the same time, panning is enabled by shifting the image laterally in response to head rotation—mimicking motion across a virtual, front-facing plane. The image remains fixed in orientation, enhancing predictability during navigation.
The Tilt condition builds on Parallel by coupling the orientation of the image plane to the user’s head pose. The zooming behavior remains identical, but the panning dynamically reorients the image to maintain perpendicular alignment with the user’s view. This technique compensates for angular displacement and maintains a consistent visual relationship between user and content, potentially reducing rotational distortion during exploration.
4 Application Scenarios
Although the primary evaluation of HeadZoom focused on a controlled image search task, the system was designed with broader use cases in mind. Head-based interaction is particularly valuable in contexts where traditional input methods are impractical or unavailable. To demonstrate the system’s generalizability, we explored how HeadZoom could be applied across several domains requiring precise, hands-free navigation of two-dimensional content.
4.1 Applications in Specific Domains
To ensure the relevance and applicability of our proposed interaction techniques, we considered potential real-world applications where efficient and natural zooming mechanisms are critical. Our selection of application areas was guided by the need for high navigation precision, frequent zoom interactions, and varying user constraints, such as the availability of hands-free control. These application areas include the following:
Geographic Information Systems (GIS): GIS interfaces rely on frequent panning and zooming. Traditional GIS software employs mouse-based interaction, while mobile applications use multitouch gestures [57]. Researchers have also explored voice- and gesture-based interactions [15].
Medical Imaging: Radiologists and pathologists navigate large and detailed images, and traditional input devices such as mice can be inefficient. Alternatives such as head tracking, gesture control, and voice commands have been evaluated for image navigation [8, 54].
Digital Photography: Photo browsing is typically based on scrolling grids and pinch-to-zoom gestures. Some researchers investigated alternative approaches such as PhotoMesa, a zoomable image browser [3]. Head-directed control offers a hands-free alternative.
To explore the broader applicability of the HeadZoom system, we extended the experimental setup to include three domain-specific use cases: GIS, medical imaging, and digital photography. These scenarios demonstrate the system’s versatility across diverse tasks and user groups, moving beyond our controlled search experiment.
Geographic Information Systems (Google Maps): The application integrates dynamic map data rendered as 256×256 pixel tiles at various zoom levels. Users are tasked with locating iconic landmarks in globally recognized cities using only head-based input: leaning forward and backward adjusts the zoom level, while head rotations pan the map horizontally and vertically (see Figure 3.a).
However, the implementation revealed technical challenges. The core issue involved mismatched numerical precision: map coordinates are expressed in double precision, while headset tracking operates with single-precision floats. Even minimal head movements could result in abrupt positional jumps at higher zoom levels. Rapid gestures often led to unintended multi-step changes, causing the system to load new tile sets and introducing visible latency.
We implemented a ’focus zoom’ technique loosely inspired on Baudisch et al. [2]. During zoom-out motions, the map area within a circular region around the user’s gaze retains the current zoom level, while the surrounding map zooms out. The inverse is applied during zoom-in actions. This approach delays full tile updates until the user’s head stabilizes, reducing perceptual discontinuities. For panning, future refinements include lower-res zoom tiles as visual placeholders, while higher-resolution tiles load incrementally.
Medical Imaging (Chest X-Rays): HeadZoom can be adapted for radiological image inspection, using high-resolution chest X-rays obtained, for example, as 12-bit DICOM files from the MIMIC data set; cf. Figure 3.b. This use case targets professional radiologists and shifts the task paradigm from target localization to clinical evaluation, similarly to [50]. Navigation across image sets is achieved using a wireless button, while a Bluetooth mouse allows users to adjust windowing parameters critical for diagnostic visibility.
Crucially, this prototype incorporates eye tracking, allowing for the collection of gaze data throughout the task. A dedicated button toggles a heatmap overlay that visualizes gaze concentration over the image, supporting both post hoc analysis and real-time feedback.
Digital Photography (Image Browsing) The third application revisits image browsing in a consumer context. Inspired by earlier systems such as PhotoMesa [3], the application enables users to explore large photo collections using head-based zoom and pan controls. Unlike GIS and medical imaging setups, which prioritize analytical precision, this scenario emphasizes intuitive navigation and user experience. The interface presents image grids where users can seamlessly zoom in individual photos and pan across the collection cf. Figure 3.c, offering a hands-free alternative to traditional gesture-based browsing.
These three scenarios exemplify the adaptability of head-based interaction across domains with varied requirements, ranging from real-time responsiveness and visual continuity in GIS, to clinical accuracy and data logging in medical imaging, and user-centric exploration in photography.

### v2

@@ Abstract
We introduce HeadZoom, a hands-free interaction technique for navigating two-dimensional visual content using head movements. HeadZoom enables fluid zooming and panning using only real-time head tracking. It supports natural control in applications such as map exploration, radiograph inspection, and image browsing, where physical interaction is limited. We evaluated HeadZoom in a within-subjects study comparing three interaction techniques—Static, Tilt Zoom, and Parallel Zoom—across spatial, error, and subjective metrics. Parallel Zoom significantly reduced total head movement compared to Static and Tilt modes. Users reported significantly lower perceived exertion for Parallel Zoom, confirming its suitability for prolonged or precision-based tasks. By minimizing movement demands while maintaining task effectiveness, HeadZoom advances the design of head-based 2D interaction in VR and creates new opportunities for accessible hands-free systems for image exploration.
@@
Zooming and panning interactions are essential for inspecting detailed visual content in high-resolution 2D images [16]. Traditional techniques rely on hand-based input, such as a mouse scroll-wheel [28], touch-screen gestures [3, 39], 6DOF controllers [55], or shortcut keys [74]. However, these methods assume that the user’s hands are free, which may not be true. With the emergence of Mixed Reality (MR), applications have expanded into critical real-world domains of surgery [31], fieldwork [66], manufacturing [54], education [59], and accessibility [14]. Here, hands are usually occupied with primary tasks, such as gripping instruments, wearing gloves, or maintaining a sterile field [43, 1]. So using a hand or a controller or touching a surface becomes inconvenient, unsafe, or impossible.
Current hands-free interaction alternatives have clear limitations. Speech suffers from recognition latency and is unusable in noise-sensitive or silence-required environments [1]. Eye-gaze techniques relying on dwell-based selection frequently lead to unintended activations, tracking inaccuracies, and visual fatigue [67]. Therefore, there is a need for more natural, precise, and practical hands-free methods. Head movements closely reflect natural user behaviours during zooming and panning operations. People instinctively lean forward to inspect details and rotate their heads to scan a scene, suggesting that head translation and rotation can be intuitive metaphors for magnification and viewport movement. Previous research [72] investigated discrete head-gesture-based zooming, but used fixed zoom increments per gesture, forcing repetitive head motions. This resulted in user fatigue and reduced precision, and notably, the method lacked support for continuous panning operations. There remains a significant research gap regarding continuous, fine-grained, and real-time control of zoom and pan via head movements.
To address this gap, we introduce HeadZoom, a novel interaction framework that maps head translation to zoom and head yaw/pitch rotations to image translation, using adaptive control-display gain to ensure smooth operation. We developed two interaction techniques (Parallel Zoom (PZ), Tilt Zoom (TZ)) and evaluated them against a baseline (Static), which uses natural head movements without specific mapping. Our results indicate that the PZ mode significantly reduces total head movement compared to both Static and TZ. In addition, participants reported substantially lower perceived exertion when using PZ, demonstrating its suitability for extended tasks. By minimizing physical demands while maintaining task performance, HeadZoom provides a viable option for hands-free 2D interactions in MR, opening new possibilities for immersive and accessible image exploration. The paper’s main contributions include:
@@
User-Centered Design: An interface that reduces cognitive and physical effort, with efficient interaction in extended use, complemented by three application scenarios and actionable design guidelines to support future practical implementations in MR environments.
Extensive Evaluation: A user study comparing the Parallel Zoom and Tilt Zoom techniques against a Static baseline.
In the rest of the paper, we first review related work, clarifying the research gap that HeadZoom addresses and our novel contribution. Next, we describe the HeadZoom implementation, before presenting a user study to evaluate the technique. After reporting on the user study results we discuss them in Section 7. We then describe application scenarios, design guidelines, and limitations. Finally, we close with conclusions and future research directions.
@@ 2 Related Work
Navigating visual content often requires panning (moving the viewport) and zooming (adjusting scale). Zooming and positioning techniques are critical to improving navigation efficiency in different applications. Many software tools, such as image editors or maps, provide interfaces for zooming- and panning-based navigation. This section reviews existing research in these areas and highlights the role of head-based interaction in addressing current challenges.
@@ 2.1 Zooming and Panning Interfaces
Zooming and panning interfaces have been proposed for the 2D desktop GUI. You et al. [74] conducted a usability evaluation of the zoom and pan functions in web-based map interfaces, comparing different zoom and pan button layouts. Gutwin and Fedak [21] used panning and zooming interfaces to fit large 2D GUIs to smaller screens. Igarashi and Hinckley [28] automatically controlled the zoom level in proportion to the scrolling speed to help users navigate large documents efficiently and smoothly, without getting disoriented. This approach is superior to manual control, which suffers from a fast visual flow when scrolling too quickly while zoomed in. Cockburn et al. [16] provide a comprehensive review of early research on interfaces for navigating large documents in 2D desktop GUI, categorizing them into Overview+Detail, Zooming, Focus+Context, and Cue-based techniques. Standard input methods for zooming and panning in 2D GUI include mouse, touch gestures, and shortcut keys [74]. In addition to these popular techniques, researchers also investigated other input methods, such as 6DOF controllers [55] or gaze [63, 23].
The popularization of mobile devices with small screens further emphasized the need to explore zooming and panning techniques [21]. Although the pinch-to-zoom method became widely adopted in touch screen-based mobile devices [3], the limitation of occupying both hands motivated many researchers to explore a one-handed solution to control zooming and panning [39]. Researchers have developed various touch gestures [18, 33] to replace or complement the standard pinch-to-zoom gesture. At the same time, others also explored using additional detectors, such as pressure sensors combined with a touch screen to enable natural zooming and scrolling [41, 65], or using a gyro sensor or buttons built into the device [20].
Zooming and panning interfaces have also been explored in AR and VR. Handheld controllers [13] or hand gestures and gaze input [58, 50, 40] are used to control zoom and pan in VR / AR, or touch-screen-based interaction [45, 12]. Lee et al. [34] proposed an automatic control of the zoom level based on the distance between the user’s view and the target object, increasing the zoom level as the AR camera approaches the tracking target, enabling users to have a closer look of the content with less tracking failure. While they linked the zoom control to the movement of the user’s view, the users held the AR camera in their hands. In comparison, our work investigates how head movement, as one of the hands-free input methods, could be useful for controlling zooming and panning to explore large, high-resolution images in VR.
@@ 2.2 Hands-Free Input Methods
Eye Tracking. Eye tracking technology has been actively researched in the context of human-computer interaction [38, 61]. Gaze direction is a natural way to control an interface, as the eyes instinctively focus on areas of interest, and has also been actively investigated in VR [52, 50] and AR [32, 47] interfaces. Modern eye trackers using infrared corneal reflections achieve high accuracy [38, 61]. A key challenge, however, is the Midas touch problem, which makes it difficult to distinguish intentional interactions from normal eye movements [30, 38]. Although some modern VR/AR HMDs integrate eye-tracking technology, this technology is mostly supported on high-end devices. It has not reached the status of being widely adopted as a standard feature.
Head-Based Interaction. Head tracking has been studied to control computers using natural head movements [27, 64]. Early work in human-computer interaction identified head movement as a simple and effective pointing method, often more accurate than voice or brain-computer interfaces [61]. It is also considered to be more affordable compared to eye-tracking technology. Head-based input is especially useful in assistive technology, allowing users with limited limb mobility to move a cursor or make selections by moving their head, using a webcam or infrared sensor. Systems such as Camera Mouse [7, 46] and Enable Viacam (eViacam)111https://eviacam.crea-si.com track the face of a user through a webcam, translating head movements into pointer control [10]. These systems have been used successfully by individuals with high spinal cord injuries, as head and eye movements remain intact [7, 61]. Harrison and Dey’s Lean and Zoom system [24] detected users leaning toward or away from a desktop screen, automatically magnifying the content. Users reported that this head-driven zooming felt natural and fluid, making the text easier to read and reducing discomfort.
Head tracking is a key enabler technology for modern VR/AR head-mounted displays (HMDs) to control the user’s viewpoint, inducing a sense of presence in VR [60, 70]. In addition, head tracking has been also actively used as a hands-free interaction method, such as using the head direction for pointing [2, 53, 71, 42]. Apart from using head-based pointing for selection, Rudi et al. [56] proposed using head-orientation based pointing to zoom and pan a map shown on a wearable display. Their method checked if the pointer went outside the set boundary to trigger zoom or pan. Wang et al. [68] proposed and tested dynamic control-display gain to control head-orientation based pointer. Their method used horizontal and vertical head rotation to control the direction of pointing. These prior work mainly used and manipulated the head orientation but did not investigate anteroposterior movements as in our work.
Researchers also actively investigated head gestures as an input method. Hou et al. [26] proposed a classifier to distinguish head gestures from head pointing to utilize both as hands-free interaction methods in a VR system. Yan et al. [72] investigated various head gestures identified through an elicitation study. While they did use anteroposterior movements to control zoom as in our work, their approach was to recognize the head gesture to trigger zoom in or out, compared to our approach which continually maps the head movement to the zoom level.
@@ 3 The HeadZoom Prototype
The HeadZoom prototype was developed in Unity3D (v6) for Meta Quest 3 headsets. HeadZoom is a hands-free interaction framework that provides precise, controller-free zooming and panning in VR. As shown in Figure 1, the system continuously captures the HMD position and orientation, smoothing through Kalman filters to suppress jitter while preserving responsiveness. A virtual image plane measuring 2 meters by 1 meter is dynamically placed at the user’s eye height and centred 2 meters in front of the user. The plane’s normal is aligned with the user’s horizontal forward direction (i.e., the forward vector with its vertical component set to zero), ensuring that the image plane remains stably in front of the user, regardless of head tilt.
@@
The HMD’s position vector provides real-time 3D coordinates of the user’s head in space. This vector is used to compute the user’s distance from the image plane, which modulates the zoom level; closer positions result in a zoom-in effect, while retreating causes a zoom-out. The forward vector represents the HMD directional orientation, specifically the normalized vector pointing where the user is facing. This vector is crucial for raycasting: It projects a ray from the user’s current position toward the image plane. Raycasting from the filtered head pose determines the target location for interaction. This plane-based mapping enables natural, spatially consistent simultaneous panning and zooming using only head motion. The intersection of this ray with the plane determines the panning target, allowing the system to infer where in the image the person is attending to. Both vectors are filtered to reduce noise before being passed to the technique controller. Zoom and pan values are clamped to predefined minimum and maximum limits to prevent content from becoming inaccessible or unstable. If tracking loss or abnormal data (such as sudden outliers) is detected, the system retains the last valid position until reliable data resumes.
3.1 Calibration and Control-Display Mapping
To accommodate the natural diversity in users’ body sizes, postures, and ergonomic preferences, HeadZoom incorporates a brief, user-specific calibration phase. The primary goal of this calibration is to ensure that the mapping between physical head movement and virtual zoom is both intuitive and safe, maximizing the usable range of interaction without causing discomfort or requiring unnatural motion. By capturing each user’s comfortable forward and backward lean limits, the system establishes personalized boundaries for zooming. This normalization process allows the full range of virtual zoom to be accessed within each individual’s natural movement envelope, eliminating the need for manual gain adjustment and preventing both overextension and underutilization. The virtual image is always positioned relative to the user’s initial pose, and as the user leans forward or backward, their current position is continuously normalized within their calibrated range. This ensures that, for every user, the closest and farthest zoom positions correspond precisely to their comfortable limits, with the image position clamped beyond these points to maintain stability and prevent accidental loss of content.
This individualized, parameterless mapping guarantees that all subsequent interactions remain responsive, accessible, and ergonomically appropriate, regardless of user variability. By prioritizing user comfort and control, the calibration phase underpins the system’s ability to deliver precise, hands-free navigation for a wide range of participants.
3.2 Interaction Techniques
As illustrated in Figure 2, HeadZoom supports three head-driven interaction techniques, each designed to systematically explore different mappings between head motion and image navigation, with an emphasis on spatial consistency, geometric predictability, and user control.
In the Static condition, the image remains completely stationary throughout the interaction: neither zooming nor panning is supported. All movements or rotations of the user’s head do not affect the position, size, or orientation of the image. This mode serves as a passive baseline for comparison, isolating the effects of head-driven navigation.
The Parallel condition implements a direct geometric mapping between head translation along the Z axis (forward/backward) and zoom level, using a linear function based on the user’s calibrated range. Panning is realized by raycasting from the user’s head position in the view direction onto a static image plane. This mapping ensures that as the user approaches the image, the angular range for lateral (yaw) and vertical (pitch) panning increases, providing broader coverage at close distances and finer control at greater distances. At the neutral position (2̃ m from the image), the maximum attainable yaw is and pitch is ; these values increase as the user leans forward. The image orientation remains fixed, which supports spatial predictability and minimizes disorientation during navigation. This approach leverages the geometric relationship between head position and the projected field of view, enabling efficient exploration of large images.
The Tilt condition extends Parallel by dynamically coupling the image plane’s orientation to the user’s head pose. The image rotates in yaw and pitch to always face the user, and in roll to match the user’s head tilt. The zoom mapping remains the same as Parallel. In this mode, the maximum attainable yaw and pitch at the neutral position are and , respectively, with these angles increasing as the user moves closer. Unlike Parallel, Roll is not artificially limited; the image rotates exactly with the user’s head, preserving alignment and supporting a consistent spatial relationship between the user and the content. This design reduces rotational distortion and maintains geometric coherence during navigation.
Note: In both Parallel and Tilt, the zoom effect is implemented as a dolly-in (camera movement) rather than a change in field of view, preserving perspective and spatial cues critical for visual search tasks.
3.3 Filtering
Because neck and head muscles naturally exhibit continuous micro-movements to maintain balance, our initial static Kalman filters—manually tuned via trial and error—smoothed most jitter but proved insufficient at higher zoom levels, where the motion-amplifying mapping magnifies even tiny shakes. To compensate, we made each filter’s Process Noise Covariance (Q) and Measurement Noise Covariance (R) dynamic functions of the user’s normalized head position along the calibrated lean interval. Specifically, we define so that zero maps to the maximum backwards lean, 0.5 to the neutral pose, and 1 to the maximum forward lean; Y is the filter parameter value. At runtime, the headset’s current X-value selects Q and R by sampling from two piecewise curves inscribed on the X–Y plane.
For the Tilt interaction, we hold Q constant at Y=0.01 across the entire range (). R remains flat at Y=0.0001 for backward-to-neutral positions (), then increases linearly from 0.0001 up to 0.1 as X goes from 0.5 to 1. In contrast, the Parallel interaction uses a mirrored Q curve—flat at 0.01 between X=0 and 0.5, then decreasing linearly to 0.0001 between X=0.5 and 1—while R follows the same piecewise profile as in Tilt. This position-based filtering adaptively dampens jitter at all zoom levels, eliminating image jerks without introducing perceptible lag.